### v1

# Influence-aware Causal Autoencoder Network for Node Importance Ranking in Complex Networks

Abstract
Node importance ranking is a fundamental problem in graph data analysis. Existing approaches typically rely on node features derived from either traditional centrality measures or advanced graph representation learning methods, which depend directly on the target network’s topology. However, this reliance on structural information raises privacy concerns and often leads to poor generalization across different networks. In this work, we address a key question: Can we design a node importance ranking model trained exclusively on synthetic networks that is effectively appliable to real-world networks, eliminating the need to rely on the topology of target networks and improving both practicality and generalizability? We answer this question affirmatively by proposing the Influence-aware Causal Autoencoder Network (ICAN), a novel framework that leverages causal representation learning to get robust, invariant node embeddings for cross-network ranking tasks. Firstly, ICAN introduces an influence-aware causal representation learning module within an autoencoder architecture to extract node embeddings that are causally related to node importance. Moreover, we introduce a causal ranking loss and design a unified optimization framework that jointly optimizes the reconstruction and ranking objectives, enabling mutual reinforcement between node representation learning and ranking optimization. This design allows ICAN, trained on synthetic networks, to generalize effectively across diverse real-world graphs. Extensive experiments on multiple benchmark datasets demonstrate that ICAN consistently outperforms state-of-the-art baselines in terms of both ranking accuracy and generalization capability.
I Introduction
Critical nodes in complex networks are of fundamental importance, as they play key roles in maintaining the network’s functionality, performance, stability, robustness, and dynamic behavior [1]. For example, in social networks, they facilitate targeted information dissemination and optimized resource allocation [2]. In transportation networks, they enable vulnerability analysis and robustness enhancement [3]. In biological networks, they help identify essential genes or proteins for therapeutic interventions [4]. With the growing scale and complexity of real-world networks, accurately identifying critical nodes has become an increasingly important yet challenging task [5, 6, 7, 8]. Developing methods that are both accurate and generalizable for node importance ranking is therefore of great theoretical and practical value [9, 10].
There has been significant research on node importance ranking in complex networks. Traditional topology-based approaches primarily rely on centrality-based measures, such as degree centrality [11], eigenvector centrality [12], and betweenness centrality [12, 13]. While some centrality measures are straightforward to compute (e.g., degree centrality), others (e.g., betweenness) can be computationally demanding, particularly in large-scale networks. Moreover, these methods focus on computing individual importance scores rather than modeling relative importance relations among nodes, which limits their effectiveness in ranking tasks. In addition, most measures capture node importance along a single structural dimension and are often tailored to specific research goals, which limits their generalizability across diverse network topologies. For instance, the degree-centrality-based method can identify high-degree hubs as the most influential nodes in a network. However, such an assumption overlooks the fact that influence can arise from nodes occupying structurally balanced or strategically positioned roles, even when their degrees are moderate.
In recent years, deep representation learning–based methods have become a powerful paradigm for node ranking. By automatically learning expressive node embeddings, these methods enable downstream tasks such as regression to be efficiently performed on the learned embeddings. This category includes Graph Convolutional Networks (GCNs) for local neighborhood aggregation [14], graph embedding techniques for dimensionality reduction [15, 16], and more sophisticated architectures such as graph attention networks [17] and graph contrastive learning frameworks [18] that capture nuanced structural dependencies.
Despite their strong representational power, most existing deep representation learning methods focus solely on modeling network structures—capturing low- and high-order proximities between nodes—while overlooking node importance information that is crucial for ranking tasks. For example, AGNN [16] and CGNN [19] typically follow a two-stage paradigm, where node representations are first learned independently using graph neural network techniques, followed by a prediction module for ranking or regression. As task-specific objectives are not integrated into the representation learning stage, the resulting embeddings often fail to capture task-relevant information, leading to suboptimal performance in node importance ranking.
Moreover, these deep representation learning–based approaches typically rely heavily on the topology of the target network. However, in many real-world scenarios, privacy constraints render the network a black box to users, who have no access to its complete structure [20]. Learning node embeddings directly from the explicit graph structure also limits the model’s ability to generalize across different networks, as such representations often overfit to specific structural patterns. These challenges motivate the development of node embedding methods that can improve the performance of importance ranking models without relying on the topology of target networks. In this work, we specifically investigate whether such embeddings can be learned through representation learning trained exclusively on synthetic networks and then applied across multiple diverse real-world networks.
Objectives and challenges. To address the limitations of existing approaches, we propose a novel representation learning framework for node importance ranking in complex networks. Our method draws inspiration from causal representation learning [21, 22, 23], which provides domain-invariant representations that generalize across different environments. As a result, it can be trained on synthetic networks and generalize effectively to diverse real-world graphs. Specifically, we pursue two key objectives: (i) to obtain effective node representation suitable for ranking without using the target network structure—enabling applicability in practical scenarios; and (ii) to improve ranking performance through mutual reinforcement between node representation learning and ranking optimization—ensuring more reliable and interpretable ranking results. Achieving these goals is nontrivial, as it raises two major challenges: (i) how to design an effective unsupervised representation learning strategy that produces network-invariant node embeddings, and (ii) how to ensure that the learned representations are inherently relevant and informative for node ranking.
Our solution. To address the first challenge, we design an influence-aware causal structure learning module within the autoencoder to capture robust low-dimensional embeddings that exhibit causal relevance to node importance. This mechanism enables the learned representations to be network-invariant and to generalize effectively to unseen target graphs. To tackle the second challenge, we formulate a unified objective function that jointly optimizes our proposed causal reconstruction and causal ranking losses, enabling a synergistic interaction between representation learning and ranking optimization.
Contributions. Our contributions are summarized as follows:
  • Influence-aware causal representation learning mechanism: We design an influence-aware causal representation learning mechanism for node ranking prediction and integrate it into an autoencoder framework. Specifically, we construct a node influence variable based on the information propagation process within the training network, and learn the causal relationships between node embeddings and this influence variable. This design encourages the learned representations to capture network-invariant causal signals related to node importance, enabling the model to be trained solely on synthetic networks and to generalize effectively to diverse real-world graphs.
  • Feature–task co-optimization mechanism: We propose the causal reconstruction loss and the causal ranking loss, and integrate them with a regularization term into a unified objective. This feature-task co-optimization framework jointly optimizes node representation learning and ranking prediction, ensuring that the resulting embeddings are directly aligned with the downstream ranking task.
  • Extensive empirical studies on real-world networks: We conduct experiments to validate the effectiveness of the proposed method. The results demonstrate that it improves the performance of the node ranking model on various real-world networks in terms of both accuracy and generalization. Ablation studies further highlight the contribution of the designed influence-aware causal mechanism.
The remainder of this paper is organized as follows. Problem formulation is introduced in Section II. The proposed influence-aware causal autoencoder network for node importance ranking is presented in detail in Section III. The experimental results are reported in Section IV. Some related work is outlined in Section V. Conclusions are drawn in the final section.
II Problem formulation
In this section, we define the problem of node ranking in a graph and introduce the key assumptions of the proposed model.
II-A Problem statement
Given a graph $G=\{V,E\}$, where $V$ is the set of nodes and $E$ is the set of edges. The notation $v_{i}\in V$ denotes a node in $V$, and $e_{ij}\in E$ denotes an edge from node $v_{i}$ to $v_{j}$. The graph can be represented by an $n\times n$ adjacency matrix $\bm{A}$, where $n$ denotes the number of nodes. The entries of the matrix are defined such that $\bm{A}_{ij}=1$ when $e_{ij}\in E$, and $\bm{A}_{ij}=0$ otherwise. Let $\bm{X}\in\mathbb{R}^{n\times d}$ denotes the feature matrix. For graphs without node attributes, $\bm{X}$ can be obtained by applying a node embedding method (e.g., Node2Vec [24]) to the adjacency matrix $\bm{A}$.
The node importance ranking is a critical problem in network analysis. Given the graph $G=\{V,E\}$, our task is to assign a score $s_{i}$ to each node $v_{i}\in V$, forming a score vector $\bm{S}=(s_{1},s_{2},\dots,s_{n})$, These scores are then used to derive a node importance ranking $\bm{R}=(r_{1},r_{2},\dots,r_{n})$, where $r_{i}$ denotes the rank of node $v_{i}$. The ranking $R$ satisfies the following: $r_{i}=r_{j}$ if $s_{i}=s_{j}$, and $r_{i}<r_{j}$ if $s_{i}<s_{j}$, meaning that a higher score corresponds to a higher (i.e., numerically lower) rank. Unlike methods that emphasize the absolute magnitude of scores, our approach focuses on preserving the relative ordering of nodes consistent with their underlying importance, making ranking consistency the primary objective.
II-B Assumptions
Our proposed method is designed to learn a robust low-dimensional node representation for predicting node importance ranking in complex networks. It is worth noting that our approach is proposed based on the following fundamental assumptions.
Assumption 1.
(Causal graph [25, 26]) For Bayesian Network $U,\mathbb{G},P>$, where $U$ denotes the set of variables, $\mathbb{G}$ is a Directed Acyclic Graph (DAG) on $U$ and $P$ denotes the probability distribution of $U$, then Bayesian Network $U,\mathbb{G},P>$ can be used to express the causal relationships between the variables in $U$. In DAG $\mathbb{G}$, for a pair of directly connected parent-child variables, the parent variable is the direct cause of the child variable, and the child variable is the direct outcome of the parent variable. We assume that the causality between variables can be expressed by $\mathbb{G}$ as a causal graph.
Assumption 2.
(Causal Markov [27, 28]) A variable $X$ is independent of every other variable (except X’s effects) conditional on all of its direct causes.
Assumption 3.
(Faithfulness [25, 29]) Every conditional independence that holds in the distribution $P$ is entailed by the Markov condition applied to $\mathbb{G}$. Faithfulness ensures that all conditional independencies in the data correspond to missing edges in the causal graph.
Figure 1: The framework of ICAN.
III Proposed method
In this section, we first show the overall framework of the proposed method, the Influence-aware Causal Autoencoder Network for node importance ranking (ICAN), and then describe its components in detail.
III-A Overview of the method
The framework of ICAN is illustrated in Fig. 1. ICAN consists of two core modules: the causal representation learning module and the causal ranking prediction module. The causal representation learning module aims to learn low-dimensional node embeddings that are causally related to node importance. Under Assumption 1, the causality among node embeddings can be described by a DAG $\mathbb{G}$. Upon obtaining the $p$-dimensional $(p\leq d)$ feature representation, the node influence score variable $\bm{Y}$, which can be used to characterize the node importance, is introduced as an additional node to form a new hidden layer. The adjacency matrix $\bm{W}=\{w_{ij}\}_{(p+1)\times(p+1)}$ associated with $\mathbb{G}$, which encodes the causal relationships among the low-dimensional node embeddings $\bm{H}^{(l)}$, is then integrated into the autoencoder framework to enable message passing under causal mechanisms. By optimizing $\bm{W}$, ICAN can learn optimal representations that are causally linked to node importance. The learned latent representations are then fed into the causal ranking prediction module, where the Markov Blanket (MB) of the node importance score variable is leveraged to predict the ranking. In this way, ICAN can produce robust, network-invariant node embeddings, enabling cross-network importance ranking predictions. In the following, we provide a detailed description of the ICAN framework.
III-B Causal representation learning module
The goal of the causal representation learning module is to infer causal relationships among low-dimensional node embeddings. In general, this module consists of two main components: an encoder and a decoder. We first generate a $d$-dimensional feature matrix $\bm{X}$ from the graph adjacency matrix $\bm{A}$ using Node2vec. The encoder, typically a Graph Convolutional Network (GCN), then takes both $\bm{X}$ and $\bm{A}$ as input. It learns to map this input into a $p$-dimensional low-dimensional representation. The decoder then uses this learned representation to reconstruct the original adjacency matrix. This integrated process of encoding and decoding can be summarized as follows:
Encoder:
${\bm{H}^{(0)}=\mathrm{ReLU}(\widetilde{\bm{A}}\bm{X}\bm{w}_{1}^{(0)}+\bm{b}_{1}^{(0)});}$ (1)
${\bm{H}^{(i)}}=\mathrm{ReLU}(\widetilde{\bm{A}}{\bm{H}^{(i-1)}}\bm{w}_{1}^{(i)}+\bm{b}_{1}^{(i)}),i=1,\cdots,m-1;$ (2)
${\bm{H}^{(m)}}=[{\bm{H}^{(m-1)}},{\bm{Y}}];$ (3)
${\bm{H}}^{(i)}=\mathrm{ReLU}(\widetilde{\bm{A}}{\bm{H}^{(i-1)}}\bm{w}_{1}^{(i)}+\bm{b}_{1}^{(i)}),i=m+1,\cdots,l.$ (4)
Decoder:
${\bm{\varPhi}^{(0)}=\bm{H}^{(l)}\bm{W};}$ (5)
${\bm{\varPhi}^{(i)}}=\mathrm{ReLU}({\widetilde{\bm{A}}\bm{\varPhi}^{(i-1)}}\bm{w}_{2}^{(i)}+\bm{b}_{2}^{(i)}),i={1},\cdots,m-1;$ (6)
${\bm{\varPhi}^{(m)}}={\bm{\varPhi}^{(m-1)}}[:,1:({p}-1)];$ (7)
${\bm{\varPhi}^{(l)}}=\mathrm{Sigmoid}(\bm{\varPhi}^{(m)}(\bm{\varPhi}^{(m)})^{T}).$ (8)
Here, $\widetilde{\bm{A}}=\bm{D}^{-\frac{1}{2}}(\bm{A}+\bm{I})\bm{D}^{-\frac{1}{2}}$ and $\bm{D}$ is the degree matrix. $l$ denotes the number of hidden layers. M[:,1:(p1)]M[:,1:({p}-1)] denotes the matrix composed of the first p1p-1 columns. Let $\bm{\varPhi}^{(l)}=\hat{\bm{A}}$ be the reconstructed adjacency matrix. $\bm{w}_{1}^{(i)}$ and $\bm{b}_{1}^{(i)}$ denote the weight matrix and bias vector on the $i^{th}$ encoding layer, respectively. Similarly, $\bm{w}_{2}^{(i)}$ and $\bm{b}_{2}^{(i)}$ denote the weight matrix and bias vector on the $i^{th}$ decoding layer, respectively. $\bm{H}^{(m-1)}\in\mathbb{R}^{n\times p}$ in the encoding process can be seen as the low-dimensional representations.
Influence-aware causal representation learning mechanism. Based on Assumption 1, learning the causal relationships among variables can be formulated as learning the matrix $\bm{W}$ associated with the DAG $\mathbb{G}$. To obtain node embeddings that are causally related to node importance, we introduce a pseudo-label variable $\bm{Y}$, termed the node influence score, which characterizes node importance based on the SIR model—a classical epidemic framework describing a spreading process where each node can be in one of three states: susceptible (S), infected (I), or recovered (R). At each time step, an infected node infects its susceptible neighbors with probability $\gamma$ and recovers with probability $\delta$, after which it cannot be infected again. The process continues until no infected nodes remain. This model captures the key dynamics of contagion and recovery, which are also relevant to information diffusion and network resilience. Therefore, it provides a reliable basis for evaluating the spreading ability of nodes and for training deep models to understand real-world network dynamics. In our experiments, each node $v$ is initialized as the only infected node, while all others are susceptible. The final fraction of infected and recovered nodes when the process stabilizes is denoted as $F_{v}$, representing the spreading influence of node $v$. The average value of $F_{v}$ over 100 simulations is used as the influence score, denoted by $\bm{Y}_{v}$. Then we run the SIR model for each node in the network to obtain the influence score variable $\bm{Y}$.
To identify the causal relationship between the node feature variables and the node influence score variable, the hidden layer must include the information of $\bm{Y}$ in the training process. Thus, starting with low-dimensional feature representation $\bm{H}^{(m-1)}$, ICAN incorporates $\bm{Y}$ as an additional node to construct a new hidden layer $H^{(m)}$ (Eq. (3)). Subsequently, $\bm{Y}$ is removed after the $(m-1)^{th}$ decoding layer (Eq. (7)).
The causal reconstruction loss $\mathcal{L}_{1}$ can be defined as
$\mathcal{L}_{1}=-\sum_{i,j}\left[\bm{A}_{ij}\log\hat{\bm{A}}_{ij}+(1-\bm{A}_{ij})\log(1-\hat{\bm{A}}_{ij})\right].$ (9)
Based on the above computational process, the reconstructed adjacency matrix $\hat{A}$ can be represented as follows:
$\hat{\bm{A}}=\sigma(\bm{\varPhi}^{(m)}(\bm{\varPhi}^{(m)})^{T}),~~\bm{\varPhi}^{(m)}=g_{2}(g_{1}(\bm{X},\bm{A})\bm{W}),$ (10)
where the adjacent matrix $\bm{W}$ is required to satisfy the constraint in Eq. (18). $g_{1}$ and $g_{2}$ can be regarded as two nonlinear functions used to describe the encoding and decoding process respectively. Therefore, it is noted here that the causal representation learning module can be seen as a graph autoencoder.
Based on the causal graph $\mathbb{G}$ or its corresponding adjacency matrix $\bm{W}$, the Markov Blanket $\text{MB}(\bm{Y})$ of the node influence score variable $\bm{Y}$ can be identified under the causal Markov assumption (Assumption 2) and the faithfulness assumption (Assumption 3). This MB($\bm{Y}$) constitutes the set encompassing the parent, child, and spouse nodes of the node influence score variable $\bm{Y}$. The robust causal relationships between the node influence score variable $\bm{Y}$ and the features within its Markov Blanket, $\text{MB}(\bm{Y})$—designated as the causal feature subset—reflect an invariant influence propagation mechanism across networks. As such, utilizing this invariant subset for node-level prediction tasks enhances model performance and generalization capability on unseen networks.
III-C Causal ranking prediction module
Next, we develop the causal ranking prediction module to assign a score to each node in the graph. The causal feature set $\text{MB}(\bm{Y})$ is utilized as the input to the ranking prediction module. The hidden layer of the rank prediction module consists of two layers of GNN and one layer of the FC layer, as expressed by the following equations.
${\bm{E}^{(0)}=\text{MB}(\bm{Y});}$ (11)
${\bm{E}^{(i)}}=\mathrm{Sigmoid}\left(\bm{A}\bm{E}^{(i)}\bm{w}_{3}^{(i)}+\bm{b}_{3}^{(i)}\right),i=1,\cdots,t;$ (12)
${\bm{E}^{(t+1)}}=\bm{E}^{(t)}\bm{w}_{3}^{(t+1)}+\bm{b}_{3}^{(t+1)}.$ (13)
Here, $\bm{A}$ represents the adjacency matrix of the graph, and $\bm{w}_{3}^{(i)}$ and $\bm{b}_{3}^{(i)}$ are the weight matrix and bias matrix respectively.
Task-aware ranking loss function. We introduce a new ranking loss function, CausalListMLE, which can be viewed as an extension of ListMLE [30] for optimizing the overall ranking prediction module. This list-wise ranking loss directly maximizes the likelihood of the entire ranked permutation, rather than focusing on individual scores as in point-wise losses like MSE. By modeling the ranking distribution, CausalListMLE focuses on preserving the correct relative order among all items, and typically achieves better performance in ranking-based learning tasks. The causal ranking loss $\mathcal{L}_{2}$ is defined as
$\mathcal{L}_{2}=-\log P(\pi_{y}\mid f_{2}(f_{1}(\mathsf{MB}(\bm{Y})))),$ (14)
where the node influence score $\bm{Y}$ is generated via the SIR model, and $\pi_{y}$ represents the corresponding ranking results. $f_{1}$ and $f_{2}$ denote the GNN layer and FC layer, respectively. The probability $P(\pi_{y}\mid f_{2}(f_{1}(\mathsf{MB}(\bm{Y}))))$ follows the Plackett–Luce model, expressed as
$P(\pi\mid x)=\prod_{i=1}^{n}\frac{\exp(x_{\pi(i)})}{\sum_{k=i}^{n}\exp(x_{\pi(k)})},$ (15)
where $x_{\pi(i)}$ is the predicted score of the node at position $i$ in the ranking $\pi$. By maximizing this likelihood, the causal ranking loss encourages the model to produce ranking that closely align with the influence score.
III-D The objective function
We have designed the causal reconstruction loss $\mathcal{L}_{1}$ to ensure the learned representations accurately causal dependencies underlying node importance, and the causal ranking loss $\mathcal{L}_{2}$ to optimize them for the downstream ranking task. In addition, to mitigate model overfitting and prevent performance degradation, we design a regularization term $\mathcal{L}_{3}$ on the weight parameters to the objective function $\mathcal{L}$:
[…]
Feature–task co-optimization mechanism. According to the previous defined losses in Eqs. (9), (14) and (16), the objective function of ICAN can be defined as
$\mathcal{L}=-\frac{\lambda_{1}}{N}\sum_{i,j}\left[\bm{A}_{ij}\log\hat{\bm{A}}_{ij}+(1-\bm{A}_{ij})\log(1-\hat{\bm{A}}_{ij})\right]$
$-\lambda_{2}\log P(\pi_{y}\mid f_{2}(f_{1}(\text{MB}(\bm{Y}))))$
[…]
To ensure that the learned $\bm{W}$ forms a DAG, this objective must satisfy the acyclicity constraint [31]:
$\mathrm{tr}\!\left(e^{\bm{W}\odot\bm{W}}\right)-(p+1)=0,$ (18)
where $\mathrm{tr}(\cdot)$ denotes the trace operator, $e^{\bm{W}}$ represents the matrix exponential, and $\odot$ denotes the Hadamard product.
Lemma 1.
Let $\mathbb{G}=(U,E)$ be a graph with adjacency matrix $\bm{W}\in\mathbb{R}^{(p+1)\times(p+1)}$. Then, the $(i,j)$-th element of $\bm{W}^{k}$, denoted by $\bm{W}_{ij}^{(k)}$, represents the number of paths of length $k$ from node $v_{i}$ to node $v_{j}$, where $i,j=1,\ldots,p+1$.
Theorem 1.
The adjacency matrix $W$ corresponds to an acyclic graph $\mathbb{G}$ if and only if Eq. (18) holds.
Proof.
Let $\bm{Q}=\bm{W}\odot\bm{W}$. Clearly, $\bm{Q}\geq 0$, i.e.,
$\bm{Q}\in\mathbb{R}^{(p+1)\times(p+1)}_{+}.$
If Eq. (18) holds, then
$\mathrm{tr}(e^{\bm{Q}})-(p+1)$ $=\mathrm{tr}(\bm{I})-(p+1)+\mathrm{tr}(\bm{Q})+\frac{1}{2!}\mathrm{tr}(\bm{Q}^{2})+\cdots$
$=\mathrm{tr}(\bm{Q})+\frac{1}{2!}\mathrm{tr}(\bm{Q}^{2})+\cdots=0.$ (19)
Since all entries of $\bm{Q}$ are nonnegative, it follows that $\mathrm{tr}(\bm{Q}^{k})=0$ for all $k\geq 1$, implying $q_{ii}^{(k)}=0$ for all $i$. By Lemma 1, there are no cycles in $\mathbb{G}$; hence, $\mathbb{G}$ is acyclic.
Conversely, if $\mathbb{G}$ is acyclic, Lemma 1 implies $q_{ii}^{(k)}=0$ for all $i$ and $k$. Thus, $\mathrm{tr}(\bm{Q}^{k})=0$, and therefore
${0=\mathrm{tr}(\bm{Q})+\frac{1}{2!}\mathrm{tr}(\bm{Q}^{2})+\cdots=\mathrm{tr}(e^{\bm{Q}})-(p+1).}$ (20)
Hence, Eq. (18) holds, proving the necessity. ∎
To solve above equality-constrained problem, the augmented Lagrangian method [32, 33] is employed to convert it into an unconstrained problem. After introducing the Lagrange multiplier and the penalty term, the objective function Eq. (III-D) is reformulated as the following augmented Lagrangian function:
$\mathcal{L}=-\frac{\lambda_{1}}{N}\sum_{i,j}\left[\bm{A}_{ij}\log\hat{\bm{A}}_{ij}+(1-\bm{A}_{ij})\log(1-\hat{\bm{A}}_{ij})\right]$
$-\lambda_{2}\log P(\pi_{y}\mid f_{2}(f_{1}(\text{MB}(\bm{Y}))))$
[…]
$+\alpha h(\bm{W})+\frac{\rho}{2}|h(\bm{W})|^{2}]{\color[rgb]{0,0,1},}$ (21)
where $\alpha$ denotes the Lagrange multiplier, $\rho>0$ denotes the penalty parameter and $h(\bm{W})=\text{tr}(e^{\bm{W}\odot\bm{W}})-(p+1)$. We then have the following update rules for the adjacent matrix $\bm{W}$, $\alpha$ and $\rho$:
$\bm{W}^{(t+1)}$ $=\arg\min~\mathcal{L}(\bm{W}^{(t)},\alpha^{(t)},\rho^{(t)}),$ (22)
$\alpha^{(t+1)}$ $=\alpha^{(t)}+\rho^{(t)}h(\bm{W}^{(t+1)}),$ (23)
$\rho^{(t+1)}$ $=\begin{cases}\beta\rho^{(t)},&\mbox{if }|\bm{W}^{(t+1)})|\leq\theta|h(\bm{W}^{(t)})|,\\ \rho^{(t)},&\text{otherwise},\end{cases}$ (24)
where $t$ denotes the $t_{th}$ iteration; $\beta>1$ and $\theta<1$ are two tuning hyperparameters. The gradient descent method is applied to solve the problem Eq. (22), thereby obtaining the adjacency matrix $\bm{W}$. Following this, the Lagrange multiplier $\alpha$ and the penalty parameter $\rho$ are updated accordingly. The proposed ICAN algorithm is shown in Algorithm 1.
Algorithm 1 ICAN
0: Adjacency matrix $\bm{A}$; parameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $l$, $m$, $t$, $d$, $\beta$, $\theta$, $\delta$, $\gamma$, $T$;
0: Prediction ranking $\hat{\bm{Y}}$;
1: Initialization: Initial weight and bias parameters randomly; Let $\bm{W}=\mathrm{\bm{I}},\rho=1,\alpha=0,t=0$.
2: Acquire node feature matrix $\bm{X}$ using node embedding methods;
3:while $t\leq T$ do
4:  Update the adjacent matrix $\bm{W}^{(t+1)}$ by Eq. (22);
5:  Update $\alpha^{(t+1)}$ and $\rho^{(t+1)}$ by Eq. (23) and Eq. (24), respectively;
6:  $t\leftarrow t+1$;
7:end while
8: Extract the causal representation $\text{MB}(\bm{Y})$ by $\bm{W}$ and $\bm{H}^{(m)}$.
9: Calculate the output of ranking prediction module $\hat{\bm{Y}}$.
10:return $\hat{\bm{Y}}$
IV Experiment
In this section, experiments are performed on the real-world networks to validate the effectiveness of the proposed ICAN method in the node importance ranking problem. In the following, the experimental settings such as the used datasets, evaluation criterion, comparison methods and model parameters for the experiments are described first, followed by a detailed analysis of the results. To systematically evaluate ICAN, we define five research questions (RQs)—our “Five CANs”:
  • RQ1: Can ICAN learn features that are causally relevant to node importance from synthetic networks—without direct access to the target network structure—and generalize well across diverse target networks?
  • RQ2: Can the feature-task co-optimization mechanism in ICAN lead to improvement in the model’s performance?
  • RQ3: Can the causal ranking loss defined in ICAN contribute to the feature representation learning process?
  • RQ4: Can the model identify important nodes that are not simply high-degree hubs, capturing more nuanced influence patterns across the network?
  • RQ5: Can the model achieve consistently high performance on one give target network when trained on different types of generative networks?
IV-A Description of datasets
We train the model on five synthetic graphs and evaluate it on six real-world networks to validate its effectiveness.
Training datasets. Five representative synthetic network models are used for training, including Barabási–Albert (BA) [34, 35], Extreme Homogeneous (EH) [36], Erdös-Rényi (ER) random-graph [37], Q-Snapback (QS) [38] and Random Hexagon (RH) [38]. The BA networks are generated using the preferential attachment mechanism, where new nodes are sequentially added and connected to existing nodes with a probability proportional to their current degrees, leading to a scale-free degree distribution. The EH networks are obtained by performing additional random edge rectifications on ER networks to transform their degree distribution from Poisson to near-uniform. The ER random graphs are generated by connecting each pair of nodes independently with a fixed probability $p$, producing networks with a binomial degree distribution. The QS networks consist of a directed backbone chain with multiple probabilistic “snapback” edges linking newly added nodes to previously added ones within a defined range, thereby enhancing backward connectivity and robustness. The RH networks are composed of randomly connected hexagonal substructures that capture local clustering and spatial organization similar to lattice-like systems. All trained networks comprise 1,000 nodes with an average degree of 4.
Target datasets. We use the following real-world networks to test the model trained on the synthetic networks:
  • Karate [39]: A human social network, in which each node represents a member in the club and each link shows the relationship between two members.
  • Jazz [40]: A human collaboration network, in which each node represents a jazz musician and each link indicates that two musicians have performed together in the same band.
  • Email-univ [41]:An email communication network from the University Rovira i Virgili in Spain, where nodes represent users and edges indicate at least one email was exchanged.
  • USAir [42]: A directed weighted network represents the flight connections among 1,574 U.S. airports in 2010. Each node corresponds to an airport, and each directed edge indicates the existence of at least one flight from the origin to the destination airport. The edge weight quantifies the total number of flights operated on that route during the year.
  • Vidal [43]: A network represents an initial version of a proteomescale map of Human binary protein–protein interactions.
  • Email-dnc [44]: A directed unweighted network derived from the 2016 Democratic National Committee (DNC) email leak. Each node represents an individual user, and each directed edge indicates that at least one email was sent from one user to another.
The key statistical properties of the real-world networks are summed up in Table I . $n$ represents the number of nodes, $m$ is the number of edges, $\langle k\rangle$ represents the average node degree, $k_{\text{max}}$ is the maximum degree, $c$ stands for the average clustering coefficient, $AS$ indicates the degree assortativity, and $HE$ is the degree heterogeneity.
TABLE I: The statistical properties of the real-world networks.
[…]
Karate 34 78 4.5882 17 0 -0.1352 3.1064
Jazz 198 2742 27.7 100 0.5376 0.0474 3.1010
Email-univ 1133 5451 9.6 71 0.2201 0.0782 1.9421
USAir 1574 17215 21.9 314 0.5042 -0.1132 5.1303
Vidal 3023 6149 4.1 129 0.0658 -0.1256 3.7960
Email-dnc 2029 4465 4.4 404 0.1948 -0.3014 16.6476
Cora 2708 5263 3.8 168 0.2389 -0.0659 2.8035
IV-B Evaluation criterion
In this section, we introduce the evaluation metric used to measure the effectiveness of the predicted ranking results by ICAN, Kendall’s $\tau$ coefficient [45]. This non-parametric statistic is widely employed to measure the ordinal association between two ranking lists. It quantifies the similarity between two ranking orders by comparing the numbers of concordant and discordant pairs. The definition of $\tau$ is
$\tau=\frac{2(N_{c}-N_{d})}{n(n-1)}$},
where $n$ is the number of items (or nodes), $N_{c}$ is the number of concordant pairs, $N_{d}$ is the number of discordant pairs, and $\tfrac{1}{2}n(n-1)$ is the total number of possible item pairs. For two rankings $\{(x_{i},y_{i})\}$ and $\{(x_{j},y_{j})\}$, a pair $(i,j)$ is considered concordant if $(x_{i}-x_{j})(y_{i}-y_{j})>0$, discordant if $(x_{i}-x_{j})(y_{i}-y_{j})<0$, and neither if $x_{i}=x_{j}$ or $y_{i}=y_{j}$. The value of Kendall’s $\tau$ lies between 1-1 and $1$: a value of $1$ indicates complete agreement between rankings, 1-1 indicates complete disagreement, and 0 indicates no correlation. This metric is particularly useful for evaluating graph-based ranking models, as it reflects how well the predicted node ranking preserves the ground-truth ordering.
IV-C Comparative methods
To demonstrate the advantages of the ICAN method in node ranking, we implement several state-of-the-art baselines. A brief summary of these methods is provided below.
  • Degree Centrality (DC) [12]. It quantifies the influence of a node based on the number of direct connections it has to other nodes in the graph. The degree centrality of a node u is formally defined as
    $\mathrm{DC}(u)=\frac{d(u)}{n-1},$
    where $d(u)$ is the degree of node $u$ (i.e., the number of edges connected to $u$), and $n$ is the total number of nodes in the network.
  • Betweenness Centrality (BC) [46]. It measures the importance of a node by quantifying how frequently it appears on the shortest paths between all pairs of nodes in the network.
    $\mathrm{BC}(u)=\sum_{s\neq t\neq u}\frac{g_{st}^{(u)}}{g_{st}},$
    where $g_{st}$ denotes the number of shortest paths from node $s$ to node $t$, and $g_{st}^{(u)}$ is the number of those paths that pass through node $u$.
  • Eigenvector Centrality (EC) [47]. It assigns each node a centrality score based not only on its own connectivity but also on the importance of the nodes it is connected to. The EC score $\bm{X}_{u}$ of node $u$ satisfies
    $\bm{X}_{u}=\frac{1}{\lambda}\sum_{k}\bm{A}_{uk}\bm{X}_{k},$
    which, in matrix form, becomes
    $\lambda\bm{X}=\bm{A}\bm{X},$
    where $\bm{A}$ is the adjacency matrix of the network, $\bm{X}$ is the eigenvector of centrality scores, and $\lambda$ is the largest eigenvalue of $\bm{A}$.
  • H-index (HI) [48]. The H-index of a node considers the degrees of its neighbors and reflects how many of them are themselves influential. For a node $u$ with neighbors $j_{1},j_{2},\ldots,j_{k}$, the H-index is defined as
    $\mathrm{HI}(u)=H\left(k_{j_{1}},k_{j_{2}},\ldots,k_{j_{k}}\right),$
    where $k_{j_{i}}$ is the degree of neighbor $j_{i}$, and $H(\cdot)$ is a function that returns the largest integer $x$ such that at least $x$ neighbors of node $u$ have degrees no less than $x$.
  • $K$-Shell (KS) [49]. The $K$-shell algorithm operates through an iterative pruning process. It iteratively peels away network layers to assign each node a $k$-value indicating its coreness based on residual connections. A higher $K$-shell value indicates greater node influence.
  • GNN-Bet [50]. This method enables node ranking by employing a dual-path aggregation mechanism that separately propagates features along incoming and outgoing shortest paths using row-wise modified adjacency matrices, with final ranking scores generated through multiplicative combination of the aggregated path representations.
  • GNN-Close [50]. This approach achieves node ranking through a hierarchical feature aggregation scheme that combines normal adjacency operations in the initial layer with column-wise modified matrices in subsequent layers to constrain feature propagation along shortest paths, producing final ranking scores via summation of layer-wise outputs.
  • RCNN [15]. This method leverages convolutional neural networks to extract features from node-centric subgraph structures and ranks node importance based on learned representations.
  • CGNN [19]. This method enhances RCNN by incorporating additional GNN layers, improving the capture of topological dependencies for more accurate vital node identification.
  • AGNN [16]. This method combines a GCN-based autoencoder with a GNN ranking head to jointly learn structural embeddings and predict node importance using listwise ranking loss.
IV-D Implementation details
The parameter settings for ICAN and the comparative methods are as follows.
For the ICAN method, set the number of hidden layers to be $l=5$ and each hidden layer to have 32 neurons. Let
$d=128,m=3,t=2,\beta=10,\theta=0.25.$
The learning rate is $\mu=0.001$ and the maximum number of iterations $T=10$. In addition, we set the recovery probability $\delta=1$ and the infection probability $\gamma=1.5\times\gamma_{c}$ to ensure effective spreading within the network. The infection threshold $\gamma_{c}$ is derived from mean-field theory [51] as
$\gamma_{c}=\frac{\langle k\rangle}{\langle k^{2}\rangle-\langle k\rangle},$ (25)
where $k$ denotes the node degree and $\langle\cdot\rangle$ indicates the average over all nodes.
For other comparison methods, all the parameters are set as default. For model evaluation, Kendall’s $\tau$ coefficient on the test networks is adopted as the primary evaluation metric. To reduce the impact of randomness on experimental results, both ICAN and the comparative methods are run 5 times each, with average values recorded. All experiments are conducted on a computer running Windows 10, equipped with an Intel(R) i5-10400 CPU at 2.90 GHz and 16 GB of memory.
IV-E Results on various real-world networks (RQ1)
In this subsection, we present a detailed analysis of the experimental results across various datasets, highlighting performance outcomes and key insights from each.
IV-E1 Training on BA-generated network and testing on various real-world networks
In the first experiment, we conduct experiments by training on the BA generated network and testing on different real-world networks to validate the model’s effectiveness. Table II presents the predicted Kendall’s correlation coefficients across different datasets. The experimental results comprehensively evaluate the proposed ICAN model against several baseline methods for node ranking across six real-world networks. As we can see, ICAN demonstrates superior and consistent performance. Notably, ICAN achieves the highest Kendall’s coefficient on every individual dataset. This culminates in the best overall average Kendall’s coefficient of 0.7707, securing the top rank among all compared methods. The fact that ICAN, trained on a BA synthetic network, generalizes effectively to diverse real-world networks highlights its robust cross-network predictive capability. This can be attributed to ICAN learning network-invariant node embeddings that transfer reliably across different networks.
TABLE II: Kendall’s coefficient of various methods(trained on BA network) on different real-world test networks.
Methods Karate Jazz Email-univ USAir Vidal Email-dnc Average Rank
DC 0.6749 0.7903 0.6236 0.6119 0.4648 0.5613 0.6211 6
BC 0.5685 0.4643 0.2374 0.4407 0.5876 0.4032 0.4503 11
EC 0.8069 0.8212 0.2534 0.6554 0.2452 0.5503 0.5554 7
HI 0.6654 0.8271 0.6854 0.6284 0.7568 0.5672 0.6884 4
KS 0.6421 0.7713 0.6828 0.6369 0.7619 0.4988 0.6656 5
GNN-Bet 0.6530 0.6789 0.3645 0.5350 0.2866 0.3662 0.4807 9
GNN-Close 0.5352 0.6551 0.3629 0.5818 0.2429 0.4442 0.4704 10
RCNN 0.7528 0.8455 0.7522 0.6709 0.9035 0.5432 0.7447 2
CGNN 0.7058 0.7911 0.6488 0.5835 0.9442 0.5649 0.7064 3
AGNN 0.6505 0.7584 0.3100 0.6429 0.4108 0.4946 0.5445 8
ICAN 0.8090 0.8504 0.7625 0.6758 0.9524 0.5740 0.7707 1
IV-E2 Training on ER-generated network and Testing on various real-world networks
Next, we conduct experiments by training on the ER generated network and testing on different real-world networks. Table III presents the Kendall’s $\tau$ coefficients of various methods evaluated across six real-world networks. As shown, ICAN attains the highest Kendall’s $\tau$ values on four out of six datasets, with an average coefficient of 0.7732, ranking first overall. This clearly demonstrates its strong generalization ability when trained on the ER-generated network and tested on structurally diverse real-world networks.
TABLE III: Kendall’s coefficient of various methods (trained on ER network) on different real-world test networks.
Methods Karate Jazz Email-univ USAir Vidal Email-dnc Average Rank
DC 0.6749 0.7903 0.6236 0.6119 0.4648 0.5613 0.6211 6
BC 0.5685 0.4643 0.2374 0.4407 0.5876 0.4032 0.4503 11
EC 0.8069 0.8212 0.2534 0.6554 0.2452 0.5503 0.5554 7
HI 0.6654 0.8271 0.6854 0.6284 0.7568 0.5672 0.6884 3
KS 0.6421 0.7713 0.6828 0.6369 0.7619 0.4988 0.6656 5
GNN-Bet 0.6137 0.6394 0.3744 0.5565 0.4125 0.4566 0.5089 9
GNN-Close 0.6887 0.2779 0.4550 0.5470 0.4079 0.4122 0.4648 10
RCNN 0.7094 0.8014 0.8141 0.6823 0.9119 0.5490 0.7447 2
CGNN 0.6749 0.6214 0.6323 0.5835 0.9442 0.5649 0.6702 4
AGNN 0.7622 0.4889 0.2950 0.6829 0.4008 0.4534 0.5139 8
ICAN 0.8092 0.8460 0.8228 0.6545 0.9524 0.5540 0.7732 1
IV-F Ablation study
Figure 2: Ablation study results.
IV-F1 Ablation study for the influence-aware causal representation learning mechanism (RQ1)
As shown in the equations of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, both are associated with $\bm{W}$, through which the node embeddings are learned to be causally related to the node influence scores. This indicates that the causal reconstruction loss and the causal ranking loss jointly contribute to the process of influence-aware causal representation learning. To emphasize the advantages of this influence-aware causal mechanism in ICAN, we conduct an ablation study by setting $\lambda_{1}=0$ and fixing $\bm{W}$ as an identity matrix. In this setup, the Markov blanket of the node influence score variable in the causal reconstruction loss is replaced by the low-dimensional representation directly obtained from the autoencoder. This variant can be regarded as a classical autoencoder applied to the ranking task and is denoted as ICAN_w/o_Cau. The comparison results between ICAN and ICAN_w/o_Cau across different networks are shown in Fig. 2(a). A noticeable performance decline of ICAN_w/o_Cau is observed in most node ranking tasks, highlighting the critical importance of the causal mechanism integrated into the autoencoder model.
IV-F2 Ablation study for the feature–task co-optimization mechanism (RQ2)
In this study, we integrate the causal ranking loss and causal reconstruction loss within a unified framework to enable causal feature extraction and ranking tasks. To evaluate the effectiveness of the proposed end-to-end feature–task co-optimization mechanism, we conduct an ablation study by transforming the integrated model into a two-stage implementation, in which the feature extraction and ranking processes are decoupled, thereby removing the synergistic optimization effect. This variant is referred to as ICAN_w/o_Syner. In the first stage, the causal reconstruction loss and a regularization term are employed as the loss function to optimize the causal representation learning module, yielding low-dimensional latent causal representations that comprehensively encapsulate the topological properties of the network. During the second stage, the learned causal representations are fed into the ranking prediction module, which is optimized using the causal ranking loss along with a regularization term to generate the predicted node influence score. These predictions are then compared against the node influence score derived from the SIR model, with Kendall’s $\tau$ coefficient used as the evaluation metric. The two-stage model is trained on the BA synthetic network and evaluated on six real-world networks. The comparison results between ICAN and ICAN_w/o_Syner on different networks are presented in Fig. 2(b). Performance decline in ICAN_w/o_Syner is observed in most of the node ranking tasks. These results demonstrate that ICAN facilitates mutual reinforcement between causal representation learning and ranking prediction, resulting in more accurate ranking outcomes.
IV-F3 Ablation study for the causal ranking loss (RQ3)
To validate the efficacy of the proposed the causal ranking loss, we conduct ablation experiments, replacing the CausalListMLE with the ListMLE loss for the node importance ranking , defined as ICAN_w/o_CausalRank, and additionally replacing the CausalListMLE with a MSE loss for the node importance score regression, denoted as ICAN_w/o_Rank. As illustrated in Fig. 2(c), the ICAN model consistently and significantly outperforms the ICAN_w/o_CausalRank and ICAN_w/o_Rank variants on all datasets. The superiority of the CausalListMLE loss stems from its inherent ability to capture ordinal relationships within ranked lists, which is more aligned with the nature of the node importance ranking objective. Overall, these findings conclusively demonstrate that integrating the causal ranking loss loss function is indispensable for achieving high-ranking accuracy, thereby affirming the design rationale behind our task-aware mechanism.
IV-G Degree distribution visualization (RQ4)
In this section, we analyze the degree distribution of nodes identified as important according to the predicted ranking. Figs. 3 and 4 compare the degree distribution of the three networks and the corresponding influence scores of the top 10% nodes identified by ICAN, DC, and CGNN. Each subfigure combines a blue histogram representing the network’s degree distribution (left $y$-axis, relative frequency) with red scatter points indicating the influence scores of individual nodes (right $y$-axis, node influence score).
A clear methodological contrast emerges. As expected for the degree–centrality-based approach, the results for DC in Fig. 3(b) and 4(b) exhibit a strong positive correlation: nodes with the highest influence scores coincide with those of the largest degrees, concentrating red points at the high-degree end of the spectrum. In contrast, ICAN and CGNN reveal markedly different patterns. High-scoring nodes appear across a broader degree range, with a noticeable concentration in the moderate-degree region. This suggests that these methods attribute high influence to nodes that are not necessarily the network’s primary hubs.
This divergence highlights a key insight: ICAN captures influence patterns that go beyond simple degree centrality. By identifying influential nodes through a more nuanced, potentially propagation-aware representation, ICAN emphasizes structurally balanced nodes that may serve as more efficient and cost-effective targets for network interventions.
Figure 3: Degree distribution of Jazz network and top $10\%$ influential nodes.
Figure 4: Degree distribution of USAir network and top $10\%$ influential nodes.
IV-H Experiments with various training networks (RQ5)
To examine the dependency of the proposed model on training networks, we conduct experiments on three real-world networks—Karate, Jazz, and Email-univ—each trained under five distinct generative graph models (BA, ER, EH, QS, and RH). As shown in Tables IV,V and VI, ICAN consistently achieves the highest Kendall’s coefficient across all generated training graphs, demonstrating its robust transferability from synthetic to real-world networks. Specifically, on the Karate network, ICAN attains an average Kendall’s coefficient of approximately 0.78, significantly surpassing all baseline methods. On the Jazz network, ICAN maintains robust and stable performance with average coefficients above 0.83, whereas competing methods exhibit substantial fluctuations across different generative networks. Similarly, on the Email-univ network, ICAN achieves the highest and most consistent results. These results collectively verify that ICAN exhibits superior flexibility and adaptability to variations in training graph structures compared with other existing approaches.
From the results, we observe that the performance of ICAN exhibits slight variations when trained on different networks for a given target network. This behavior may stem from the structural similarities between the training and target networks. In future work, we plan to conduct a more systematic investigation of this phenomenon and establish a theoretical framework for selecting or generating suitable training networks for specific target networks.
TABLE IV: Kendall’s coefficient of various methods on Karate network.
Methods BA ER EH QS RH
GNN-Bet 0.6530 0.6137 0.6280 0.6280 0.5602
GNN-Close 0.5352 0.6887 0.6780 0.6244 0.6958
RCNN 0.7528 0.7094 0.7130 0.7094 0.7058
CGNN 0.7058 0.6749 0.6749 0.6737 0.6737
AGNN 0.6505 0.7694 0.7586 0.6865 0.5828
ICAN 0.8090 0.8092 0.7910 0.7622 0.7442
TABLE V: Kendall’s coefficient of various methods on Jazz network.
Methods BA ER EH QS RH
GNN-Bet 0.6789 0.6394 0.6711 0.6485 0.7178
GNN-Close 0.6551 0.2779 0.4140 0.5574 0.5806
RCNN 0.8455 0.8014 0.8012 0.8156 0.8212
CGNN 0.7911 0.6214 0.6214 0.6219 0.6219
AGNN 0.7584 0.3610 0.7338 0.2400 0.2926
ICAN 0.8400 0.8504 0.8303 0.8299 0.8321
TABLE VI: Kendall’s coefficient of various methods on Email-univ network.
Methods BA ER EH QS RH
GNN-Bet 0.3645 0.3744 0.3646 0.4024 0.3614
GNN-Close 0.3629 0.4550 0.4662 0.5089 0.4616
RCNN 0.7522 0.8141 0.8149 0.7878 0.7845
CGNN 0.6488 0.6323 0.6325 0.6323 0.6326
AGNN 0.3100 0.3485 0.2995 0.2800 0.2889
ICAN 0.7625 0.8219 0.8222 0.7838 0.7737
IV-I Case study for attribute graphs
Previously, we conducted experiments on six real-world networks without attributes, where node features are generated by Node2Vec. In this case study, we show that ICAN can be applied to an attributed network. The feature matrix can be constructed from node attributes. Cora [52] is a citation network where nodes represent scientific publications and edges represent directed citation links between them. Each node feature in the Cora dataset is a 1,433-dimensional binary bag-of-words vector representing the presence or absence of specific terms in the paper’s content. Our model is trained on BA synthetic networks (using Node2Vec for feature generation) and evaluated on the attributed Cora real-world network. The experimental results, as shown in Fig. 6, demonstrate that our method achieves superior ranking performance compared to all baseline approaches. This indicates that ICAN effectively generalizes across different types of networks and is adept at integrating both topological and attribute information for improved representation learning.
Figure 5: Parameters sensitivity study on all the real-world networks for ICAN.
Figure 6: Kendall’s $\tau$ of different methods on Cora dataset.
IV-J Parameter sensitivity analysis
To investigate the sensitivity to the balanced parameters $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$, we conduct a parameter sensitivity analysis of the models trained on the BA network, evaluating their performance on various real-world networks.
Let $\lambda_{2}=\lambda_{3}=1$, the Kendall’s coefficient with different values of $\lambda_{1}$ are shown in Fig. 5(a). Similarly, setting the values of the other two parameters to be the same, the results with difference values of $\lambda_{2}$ and $\lambda_{3}$ are displayed in Fig. 5(b) and Fig. 5(c) respectively. As shown in the figures, ICAN is sensitive to the parameters $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$. Specifically, for $\lambda_{1}$, the optimal value range is $[0.3,0.5]$ for Karate, and $[0.5,1.5]$ for the other networks. As for $\lambda_{2}$, the range of $[0.1,1.5]$ is optimal for the Vidal, as well as the Email-univ and Email-dnc networks, $[0.1,0.3)$ for Jazz, $[1,1.5)$ for USAir and Karate. Regarding $\lambda_{3}$, the optimal value range is $(0.3,0.5]$ for Jazz, $[1,1.2)$ for USAir,and $[0.5,1.5]$ for the other networks. Based on these findings, in the implementation of ICAN, the parameters $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ can be set to (0.5, 1, 1.5) for the Karate network, (1, 0.1, 0.5) for the Jazz network, (0.5, 1, 1) for the Email-univ network, (1, 1, 1) for the other networks, respectively.
V Related work
Our framework is a node-ranking method built on causal representation learning. Below, we review the two research areas most closely related to our model.
Node ranking methods: Previous works for node importance ranking can be categorized into two main groups: traditional topological-based methods and deep representation learning-based methods. Over the past decades, numerous topology-based node ranking methods have been developed, which can be broadly categorized into neighborhood-based, eigenvector-based, and path-based approaches. Neighborhood-based methods, such as degree centrality [53], H-index [54], and $K$-shell decomposition [49], assess node importance using local structural information. In contrast, eigenvector-based methods like eigenvector centrality [12, 55] and PageRank [56] capture global influence propagation. Path-based methods, including betweenness centrality [46] and eccentricity centrality [57], evaluate nodes based on their positions and roles in network paths. Apart from non-deep techniques, GCN based deep represeantion learning models have shown their superiority over most of existing methods. Zhao et al. [14] proposed the InfGCN algorithm, which takes neighbor graphs and classic structural features as the input into a graph convolutional network for learning nodes’ representations, and then feeds the representations into a classifier layer. Keikha et al. [58] proposed DeepIM that employs network embedding via deep learning to learn node representations preserving local and global network structures for influence maximization, particularly addressing the context of interconnected social networks. Zhang et al. [19] combined convolutional neural networks (CNN) with GNNs in their CGNN algorithm, which simplifies feature matrices and focuses on first- and second-order neighbors to label nodes via the SIR model [59] and optimize a loss function for precise identification of critical nodes. Xiong et al. [16] introduced the AGNN algorithm, which combines autoencoders with GNNs to create topological feature embeddings using GCNs and optimizes ranking predictions via listMLE. Ahmad et al. [60] proposed frameworks (LCNN) that merge CNNs with local node representations and multi-scale metrics. Yu et al. [15] developed RCNN, a method that integrates GCNs with CNN-based adjacency features for efficient critical node detection. Munikoti et al. [61] proposed ILGR, an inductive Graph Neural Network framework that leverages local sub-graph embeddings and a ranking loss to efficiently identify critical nodes and links in large complex networks based on graph robustness metrics. Huang et al. [62] introduced HIVEN, a GNN-based framework designed for heterogeneous information networks (HINs).
These node-ranking methods face several limitations: they rely on fully observable target network topology, overfit to specific structural patterns—hindering cross-network generalization—and often use learning objectives misaligned with the final ranking task. In contrast, ICAN learns robust, low-dimensional embeddings without relying on the topology of target networks, enabling strong generalization to unseen graphs.
Causal representation learning methods: Causal relationships have the ability to capture the fundamental mechanism of data generation and are stable across various contexts [63, 22]. Learning causal representations is significant for enhancing the robustness of predictive models. Zheng et al. [31] introduced NOTEARS, a novel method that reformulates the DAG discovery problem as a continuous optimization task incorporating acyclicity constraints, enabling solution via numerical techniques. Subsequently, Ng et al. [32] expanded upon NOTEARS within a graph autoencoder framework, leveraging multilayer perceptrons (MLPs) to capture and model nonlinear structural equations, thereby advancing the capability to handle complex, non-linear relationships in causal inference. Yang et al. [64] integrated a deep autoencoder and a causal structure learning model to learn causal representations using data. Yu et al. [65] proposed DAG-GNN, a deep generative model that uses a graph neural network-based variational autoencoder to learn DAG structures from data, effectively generalizing linear structural equation models to capture nonlinear relationships and handle diverse variable types.
Most existing causal representation learning methods aim to uncover causal relationships among variables, but they are not tailored for downstream ranking tasks, making it difficult to ensure the learned representations are effective for ranking. To this end, ICAN integrates a causal ranking loss into its representation learning model, enabling it to produce robust, low-dimensional node embeddings well suited for node-importance ranking tasks.
VI Conclusions
This paper proposes an Influence-aware Causal Autoencoder Network, named ICAN, for node importance ranking in complex networks. The proposed method effectively captures the network-invariant causal relationships between node features and influence scores. This design allows ICAN to learn robust, low-dimensional node embeddings solely from synthetic networks, which can then be applied to node-ranking tasks on various real-world networks. We conduct comprehensive experiments by training ICAN on five types of synthetic networks and evaluating it on diverse benchmark networks, comparing it performance with eleven representative node-ranking methods. The results confirm that ICAN achieves state-of-the-art performance, excelling in both ranking accuracy and cross-network generalization. The results of the ablation studies confirm that both the influence-aware causal representation learning mechanism and the feature–task co-optimization strategy—based on the jointly defined causal ranking loss and causal reconstruction loss—play crucial roles in ensuring that the learned representations generalize effectively across diverse target graphs and are better suited for downstream ranking tasks, thereby substantially enhancing the model’s overall performance. Future work will focus on establishing a theoretical framework for selecting or generating suitable training networks for a given target network, as well as investigating the conditional generalization principle, for the node importance ranking task on one given network.
VII AI-Generated Content Acknowledgment
We did not use any artificial intelligence (AI)-generated content (e.g., text, figures, images, or code) in this article.